Separate 65into two parts so that 3times the larger will be 6 more than 6 times the smaller.
Answer by stanbon(57967)

Separate 65into two parts so that 3times the larger will be 6 more than 6 times the smaller.
-------
Equation:
3x = 6(65-x)+6
-----
x = 2(65-x) + 3
x = 130-2x+3
3x = 133
x = 44 1/3
-----------------------
Cheers,
Stan H.
==========================